mathslover (mathslover):

If : \(l_1 = \lim_{x \rightarrow 2} \left( x + |x| \right) \) \(l_2 = \lim_{x\rightarrow 2} (2x - |x| ) \) and \(l_3 = \lim_{x \rightarrow \frac{\pi}{2} } \left( \cfrac{ \cos x}{x - \cfrac{\pi}{2}} \right) \) Find the relationship between \(l_1\) , \(l_2\) and \(l_3\) .

mathslover (mathslover):

It might not be tough for @hartnn , @mukushla or @Miracrown but, this is the question, where I'm stuck. Its not a challenge now ..!

hartnn (hartnn):

whats stopping you from direct substitution in l1 and l2 ?

Miracrown (miracrown):

i take it as a challenge :P

OpenStudy (anonymous):

:-p

Miracrown (miracrown):

Well, I suppose if we just compute them, we can come up with any number of relationships we want.

mathslover (mathslover):

@hartnn didn't get you , sorry! Did you mean to say that here is a substitution also?

hartnn (hartnn):

\(\Large \lim \limits_{x \to a}f(x) =f(a)\)

OpenStudy (goformit100):

Refet to Tata Mc-Graw Hill Series Mathematics Book.

mathslover (mathslover):

Okay, I see. So, \(l_1 = 4\) and \(l_2 = 0\) and \(l_3 = 0/0\) I don't have to do direct substitution in l_3 , right?

Miracrown (miracrown):